A spherical balloon is deflated so that its volume is decreasing at a rate of 3 ftଷ/min. How fast is the diameter of the balloon decreasing when the radius is 2 ft? 


1
Expert's answer
2021-09-28T01:11:43-0400

"V=\frac{4}{3}\pi r^3=\frac{1}{6}\pi d^3"


"\frac{dV}{dt}=\frac{d}{dt}(\frac{4}{3}\pi r^3)=\frac{4}{3}\pi\cdot 3r^2\frac{dr}{dt}=3"


"\frac{dr}{dt}=\frac{3}{4\pi r^2}"


"\frac{dD}{dt}=\frac{3}{\pi D^2}"


"\frac{dD}{dt}(4)=\frac{3}{\pi 4^2}=\frac{3}{16\pi }=0.06" ft/min